# Partnering with Generative AI: Experimental Evaluation of Human-Led and Model-Led Interaction in Human-AI Co-Creation

### v1

@@ Abstract.
Large language models (LLMs) show strong potential to support creative tasks, but the role of the interface design is poorly understood. In particular, the effect of different modes of collaboration between humans and LLMs on co-creation outcomes is unclear. To test this, we conducted a randomized controlled experiment ($N=486$) comparing: (a) two variants of reflective, human-led modes in which the LLM elicits elaboration through suggestions or questions, against (b) a proactive, model-led mode in which the LLM independently rewrites ideas. By assessing the effects on idea quality, diversity, and perceived ownership, we found that the model-led mode substantially improved idea quality but reduced idea diversity and users’ perceived idea ownership. The reflective, human-led mode also improved idea quality, yet while preserving diversity and ownership. Our findings highlight the importance of designing interactions with generative AI systems as reflective thought partners that complement human strengths and augment creative processes.
@@
$c$s: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanitiesccs: Human-centered computing Empirical studies in HCI
Figure 1. Overview of the study procedure. (A) Participants ($N=486$) were randomly assigned to one of five conditions: four LLM variants (vanilla, model-led, suggestion, question) or a control group without support. (B) Before submitting their final ideas, all participants were instructed to refine their idea using the LLM. (C) Finally, idea quality was assessed by expert raters with domain expertise in the automotive industry; idea diversity was measured via text-embedding similarity of idea descriptions; and perceived ownership was self-reported by participants for the ideas they generated. : Study procedure with five conditions, refinement, and evaluation.
@@ 1. Introduction
While existing studies on LLMs in creative contexts have examined the creativity abilities of LLMs (see Section 2 and the overview paper in (Holzner et al., 2025)), the role of the interaction design is largely ignored. In particular, existing empirical studies focus on a simple interaction format (”one-shot“) (Holzner et al., 2025; Hubert et al., 2024; Koivisto and Grassini, 2023), where the input to the model is limited to a single-prompt, so that iterative refinements are ignored. Yet, creativity theory stipulates that creativity is not a one-off act but an ongoing, cyclical process of generating, evaluating, and refining ideas (Wallas, 1926; Guilford, 1967; Frich et al., 2019). Hence, it is natural that also LLMs for co-creation should be designed as iterative thought partners that scaffold human reflection rather than as autonomous creators (Rafner et al., 2023). Motivated by this, we stipulate that LLMs can be integrated in two main ways: (1) model-led, with the model taking initiative by rewriting or expanding ideas, or (2) human-led, with the model eliciting elaboration through questions or suggestions while keeping humans in control of the creative process. Yet, empirical evidence on how different human-led vs. model-led interaction designs and thus general HCI recommendations for such co-creativity systems are lacking. Hence, we explore how different designs may support perceived idea ownership, leverage idea quality, and preserve idea diversity in creative processes.
@@
To address this research question, we conducted a pre-registered randomized controlled experiment ($N=486$). In particular, we compared five different conditions: (1) a human-led question-mode, where the LLM asked reflective questions that should trigger human thinking; (2) a human-led suggestion-mode, where the LLM provided tips to refine the idea; (3) a model-led mode in which LLMs proactively refined participants’ ideas; (4) a vanilla LLM, with the default behavior of the LLM that was left unchanged; and (5) a control group without any LLM support. Participants were asked to generate creative ideas using an established task, namely, the Alternative Uses Test (AUT) (Guilford, 1967), which involves repurposing existing objects (i.e., finding new applications for car sensors). Afterwards, ideas were evaluated along three dimensions: idea quality, idea diversity, and perceived idea ownership.
@@
    Empirically, we provide the first large-scale controlled experiment ($N=486$) directly comparing model-led and human-led interaction modes with LLMs in a complex creativity task, which shows that model-led idea refinement improves idea quality at the cost of idea diversity and perceived idea ownership. In contrast, actively involving participants with a question-driven interaction mode produces high-quality ideas while preserving ownership and even enhancing diversity.
@@
Model capabilities for creativity LLMs using one-shot prompts or no default LLMs (like our vanilla bot) match or surpass average humans on established benchmarks for divergent thinking (e.g., Alternative Uses Test [AUT]) (Lee and Chung, 2024; Hubert et al., 2024). However, while LLM support boosts humans on average, the best humans still surpass LLMs in benchmark tests (Koivisto and Grassini, 2023). How do LLMs perform in interactive co-creation processes that reflect real-world application scenarios?
Designing systems for co-creation Research on co-creation using AI can be loosely grouped into three main streams: Artifacts/system contributions: Tailored systems have been developed as creativity support tools, often using LLMs, such as for creative problem solving, collaborative songwriting (Kim et al., 2025), group ideation tools (Shaer et al., 2024), or non-linear design processes (Zhou et al., 2024). Theoretical papers: This stream characterizes creativity processes when using such systems, finding that creativity unfolds in iterative, non-linear ways, and therefore advocates AI as a collaborative thought partner instead of a tool for simple command execution (Rafner et al., 2023; Zhou et al., 2024). Further, being exposed to AI content during such creativity processes can lead to idea fixation (Wadinambiarachchi et al., 2024). Experimental evaluations: Some design choices are studied experimentally, such as, the timing (who proposes the idea first, i.e., human or AI) (Qin et al., 2025). What are effective interaction modes, where creativity emerges in the collaboration with the LLM itself?
Tensions in human-AI co-creation Prior work shows several tensions that arise in the co-creation: Decrease in idea quality: People perform worse unaided after interacting with the AI than before (Kumar et al., 2025). Homogenization effect: Interaction with LLMs (and LLMs alone) reduces the diversity of the idea pool (Doshi and Hauser, 2024; Anderson et al., 2024b; Kadoma et al., 2024) Psychological side-effects: Co-creating with AI reduces psychological ownership (Joshi and Vogel, 2025; Draxler et al., 2024). How can we address these tensions?
@@ 2.2. Augmenting Human Creativity
In HCI, scholars have long been interested in supporting human creativity through technology, often subsumed under the umbrella of creativity support tools (CSTs) (Shneiderman, 2002). CSTs have been applied across various domains, including visual design, storytelling, and ideation platforms (Frich et al., 2019). Previous creativity research in HCI has evolved in waves with two peaks in the 1990s (individual tools) and 2000s (collaborative platforms) (Frich et al., 2019), while a third wave is seen in response to the advances of large language models, where technology might have reached a stage where it can, to some extent, exhibit creativity on its own. Here, creativity is typically assessed through standardized tests such as the Alternative Uses Test (AUT), which asks to generate as many novel and diverse uses as possible for a common object, such as a brick (Guilford, 1967). In these benchmark tests, LLMs are very powerful and can even surpass average human performance (Hubert et al., 2024). However, in more complex creative tasks, such as creating a new innovative product for a start-up or writing novels, LLMs remain less capable on their own (Holzner et al., 2025). What is still missing, however, is an experimental evaluation of whether and how collaboration between humans and LLMs can augment creativity in such complex settings.
From research in social psychology, we know that creativity emerges from social interactions in the moment, where people listen, respond, and build on one another (Sawyer, 2019). Like in jazz, one instrument starts to play and others improvise on the spot, creating a completely new tune on the spot. Similarly, in the innovation context, new products might emerge in an exchange of people with different experiences and backgrounds, like designers, computer scientists, and economists. However, this iterative and collaborative process remains underexplored in current HCI research on LLMs and creativity.
@@
Psychological theory offers ways to formalize the challenges in human collaborative settings: social loafing refers to the tendency of individuals in a group to contribute less effort compared to when they work alone (Latané et al., 1979; Karau and Williams, 1993). For example, in a brainstorming session, individuals may put less effort into generating novel ideas when working in a large group compared to working alone (Diehl and Stroebe, 1987; Ez-Zaouia and Carrillo, 2024). Especially when lacking control over an outcome, humans might be demotivated and spend fewer resources (Latané et al., 1979; Maier and Seligman, 1976). Translated into human–AI collaboration, this tendency can manifest as algorithmic loafing: rather than engaging fully, individuals reduce their own effort and adopt the system’s output with little reflection (Inuwa-Dutse et al., 2023). Applied to our creativity context, this means that, instead of actively striving for the best solution, people may disengage and settle for whatever the AI produces.
@@
As a consequence, in the context of co-creating with LLMs, it might be beneficial to keep humans engaged and in charge of the creative process. Co-creation with LLMs often reduces human effort and makes the task less demanding (Mei et al., 2025). However, research shows that the mode of collaboration affects creative performance. When people co-create with a model rather than merely edit its output, they report higher self-efficacy, which, conversely, leads to higher creative quality (McGuire et al., 2024). Consequently, effective co-creation keeps humans involved in the creative process and triggers their creativity rather than allowing mental offloading.
@@
Extending this perspective, creativity research provides support on how to foster creative processes. People produce more novel ideas when they are provided analogies from unrelated domains that are still structurally relevant (Dahl and Moreau, 2002). These trigger cognitive processes and allow for producing novel content. CSTs might support here, for example, by providing analogies that are distinct but connected (Gilon et al., 2018). Building on this insight, we argue that human–LLM interaction modes are more effective when they stimulate the human creative process rather than letting the model take the creative lead. For example, when exploring new product ideas, individuals may either become constrained by the model’s output or be inspired by novel analogies they might not have considered otherwise. Here, we distinguish between (1) model-led collaboration, where the LLM generates the core ideas and the human mainly edits or reviews, and (2) human-led collaboration, where the human drives ideation and the LLM supports through providing analogies, useful information, and supporting in elaboration.
@@ 3.2. Group Level: Idea Diversity
While co-creation with LLMs might augment creativity on an individual level, this may still reduce diversity at the group level. Ideas generated by LLMs are more similar to each other compared to ideas generated by a human crowd, implying a possible homogenization effect (Meincke et al., 2025). Also, when humans collaborate with LLMs, the homogenization effect applies, so that ideas (or other text such as stories) may become more similar to each other and thus less diverse (Doshi and Hauser, 2024; Meincke et al., 2024). A reason for this effect can be human bias, such fixation effects (Kohn and Smith, 2011; Sio et al., 2015; Crilly, 2019), which also emerge in human-to-human collaborative settings. In contrast to this, we suggest that this effect could be mitigated by interaction modes, where the humans do the main creative work and the model rather acts as a source of inspiration compared to model-led conditions:
@@ 4. Methods
To empirically assess our hypotheses, we performed a preregistered (https://aspredicted.org/nf5r-cmrt.pdf) randomized controlled experiment with $N=502$ participants. Codes and anonymized data are available via anonymized GitHub (https://anonymous.4open.science/r/CreativityExperiment-4D64). Upon acceptance, we will move both code and data to a public GitHub repository. Our study compared human-led idea generation (question-mode and suggestion-mode) and model-led idea generation, along with a vanilla LLM condition and a no-LLM control condition. Our study received IRB approval from [University Name] [IRB/Ethics Committee] (Protocol [ID], approved March 12, 2025). All participants provided their informed consent and could withdraw at any time. Data were collected and stored in accordance with local data privacy laws. Compensation was done via the Prolific platform, meeting the fair‑wage policy of the platform.
Figure 2. Screenshots of the interface of the two human-led conditions. In the question-mode, the LLM provided (A) further knowledge and (B) an analogy to inspire human creativity. After defining three alternative ideas, the user chooses one to elaborate further. In the suggestion-mode, the participants were provided three suggestions on how to make the initial idea more novel, and on how to further elaborate (C). Below the suggestion, a text box is shown where the participants can refine their initial idea based on the LLM input (D). : Interfaces of question-mode and suggestion-mode with example dialogues.

### v2

@@ Abstract.
Large language models (LLMs) show strong potential to support creative tasks, but the role of the interface design is poorly understood. In particular, the effect of different modes of collaboration between humans and LLMs on co-creation outcomes is unclear. To test this, we conducted a randomized controlled experiment ($N=486$) comparing: (a) two variants of reflective, human-led modes in which the LLM elicits elaboration through suggestions or questions, against (b) a proactive, model-led mode in which the LLM independently rewrites ideas. By assessing the effects on idea quality, diversity, and perceived ownership, we found that the model-led mode substantially improved idea quality but reduced idea diversity and users’ perceived idea ownership. The reflective, human-led mode also improved idea quality, yet while preserving diversity and ownership. We independently validated the findings in a different context ($N=640$). Our findings highlight the importance of designing interactions with generative AI systems as reflective thought partners that complement human strengths and augment creative processes.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791185isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanities
Study workflow across five experimental conditions Study workflow diagram with three labeled parts. Panel A shows five conditions: control (participant writes an initial idea only), vanilla chatbot (chat interface with a short exchange), model-led (participant provides an initial idea and the system rewrites it), suggestion-mode (system presents three suggestions), and question-mode (system asks reflective questions such as ”What if” and ”Explain”). An arrow leads to Panel B, where the participant submits a refined idea. Panel C shows evaluation of the refined ideas on three outcomes: idea diversity, idea quality, and perceived idea ownership.
Refer to caption Study procedure with five conditions, refinement, and evaluation. The figure illustrates the study procedure. Panel A shows five experimental conditions in which participants first write down an initial idea. In the control condition, they receive no support. In the vanilla condition, the initial idea is directly entered into the system. In the model-led condition, the system requires additional input from the participant. In the suggestion-mode condition, the system provides multiple suggestions. In the question-mode condition, the system asks the user questions. Panel B shows that all participants can then refine their idea before submitting it. Panel C shows the evaluation stage, where ideas are assessed for diversity, quality, and perceived ownership.
Figure 1. Overview of the study procedure of the main study. (A) Participants ($N=486$) were randomly assigned to one of five conditions: four LLM variants (vanilla, model-led, suggestion, question) or a control group without support. (B) Before submitting their final ideas, all participants were instructed to refine their idea using the LLM. (C) Finally, idea quality was assessed by expert raters with domain expertise in the automotive industry; idea diversity was measured via text-embedding similarity of idea descriptions; and perceived ownership was self-reported by participants for the ideas they generated.
@@ 1. Introduction
While existing studies on LLMs in creative contexts have examined the creativity abilities of LLMs (see Section 2 and the overview paper in (Holzner et al., 2025)), the role of the interaction design is largely ignored. In particular, existing empirical studies focus on a simple interaction format (”one-shot“) (Holzner et al., 2025; Hubert et al., 2024; Koivisto and Grassini, 2023), where the input to the model is limited to a single prompt, so that iterative refinements are ignored. Yet, creativity theory stipulates that creativity is not a one-off act but an ongoing, cyclical process of generating, evaluating, and refining ideas (Wallas, 1926; Guilford, 1967; Frich et al., 2019). Hence, it is natural that also LLMs for co-creation should be designed as iterative thought partners that scaffold human reflection rather than as autonomous creators (Rafner et al., 2023). Motivated by this, we stipulate that LLMs can be integrated in two main ways: (1) model-led, with the model taking initiative by rewriting or expanding ideas, or (2) human-led, with the model eliciting elaboration through questions or suggestions while keeping humans in control of the creative process. Yet, empirical evidence on how different human-led vs. model-led interaction designs and thus general HCI recommendations for such co-creativity systems are lacking. Hence, we explore how different designs may support perceived idea ownership, leverage idea quality, and preserve idea diversity in creative processes.
@@
To address this research question, we conducted a pre-registered randomized controlled experiment ($N=486$). In particular, we compared five different conditions: (1) a human-led question-mode, where the LLM asked reflective questions that should trigger human thinking; (2) a human-led suggestion-mode, where the LLM provided tips to refine the idea; (3) a model-led mode in which LLMs proactively refined participants’ ideas; (4) a vanilla LLM, with the default behavior of the LLM that was left unchanged; and (5) a control group without any LLM support. Participants were asked to generate creative ideas using an established task, namely, the Alternative Uses Test (AUT) (Guilford, 1967), which involves repurposing existing objects (i.e., finding new applications for car sensors). Afterwards, ideas were evaluated along three dimensions: idea quality, idea diversity, and perceived idea ownership. We independently validated our findings on LLM-enabled co-creation in the more naturalistic setting of product ideation (i.e., in a second, preregistered experiment with $N=640$ participants).
@@
    Empirically, we provide the first large-scale controlled experimental evaluation (total $N=1,126$) directly comparing model-led and human-led interaction modes with LLMs in a complex creativity task, which shows that model-led idea refinement improves idea quality at the cost of idea diversity and perceived idea ownership. In contrast, actively involving participants with a question-driven interaction mode produces high-quality ideas while preserving ownership and even enhancing diversity.
@@
Model capabilities for creativity LLMs using one-shot prompts or default LLMs (like our vanilla bot) match or surpass average humans on established benchmarks for divergent thinking (e.g., Alternative Uses Test [AUT]) (Lee and Chung, 2024; Hubert et al., 2024). However, while LLM support boosts humans on average, the best humans still surpass LLMs in benchmark tests (Koivisto and Grassini, 2023). How do LLMs perform in interactive co-creation processes that reflect real-world application scenarios?
Designing systems for co-creation Research on co-creation using AI can be loosely grouped into three main streams: Artifacts/system contributions: Tailored systems have been developed as creativity support tools, often using LLMs, such as for creative problem solving, collaborative songwriting (Kim et al., 2025), group ideation tools (Shaer et al., 2024), or non-linear design processes (Zhou et al., 2024). Theoretical papers: This line of research outlines theoretical frameworks describing how creativity unfolds in co-creation with technology. It emphasizes that creativity unfolds in iterative, non-linear ways, and therefore advocates AI as a collaborative thought partner instead of a tool for simple command execution (Rafner et al., 2023; Zhou et al., 2024; Collins et al., 2024). Further, being exposed to AI content during such creativity processes can lead to idea fixation (Wadinambiarachchi et al., 2024). Experimental evaluations: Some design choices are studied experimentally, such as, the timing (who proposes the idea first, i.e., human or AI) (Qin et al., 2025). What are effective interaction modes, where creativity emerges in the collaboration with the LLM itself?
Tensions in human-AI co-creation Prior work shows several tensions that arise in the co-creation process: Decrease in idea quality: People perform worse unaided after interacting with the AI than before (Kumar et al., 2025). Homogenization effect: Interaction with LLMs (and LLMs alone) reduces the diversity of the idea pool (Doshi and Hauser, 2024; Anderson et al., 2024b; Kadoma et al., 2024). Psychological side-effects: Co-creating with AI reduces psychological ownership (Joshi and Vogel, 2025; Draxler et al., 2024). How can we address these tensions?
@@ 2.2. Augmenting Human Creativity
In HCI, scholars have long been interested in supporting human creativity through technology, often subsumed under the umbrella of creativity support tools (CSTs) (Shneiderman, 2002). CSTs have been applied across various domains, including visual design, storytelling, and ideation platforms (Frich et al., 2019). Previous creativity research in HCI has evolved in waves with two peaks in the 1990s (individual tools) and 2000s (collaborative platforms) (Frich et al., 2019), while a third wave is seen in response to the advances of large language models, where technology might have reached a stage where it can, to some extent, exhibit creativity on its own. Here, creativity is typically assessed through standardized tests such as the Alternative Uses Test (AUT), which asks to generate as many novel and diverse uses as possible for a common object, such as a brick (Guilford, 1967). In these benchmark tests, LLMs are very powerful and can even surpass average human performance (Hubert et al., 2024). However, in more complex creative tasks, such as creating a new innovative product for a start-up or writing novels, LLMs remain less capable on their own (Holzner et al., 2025). Kumar et al. (2025) empirically show that exposure to LLM-generated ideas and strategies can increase performance during creativity benchmarks but can hinder later unaided generation. However, the assistance in their study was static and thus without interactive elements: participants viewed pre-generated lists of ideas or cues rather than engaging in a dynamic dialogue with an LLM, leaving open how interactive co-creation might shape creative performance. HCI research and theory, such as the Co-Creative Framework for Interaction (COFI) (Rezwana and Maher, 2023), emphasize the relevance of the specific interaction design to create effective co-creation. Still, there is no experimental evaluation of whether and how interactive collaboration between humans and LLMs can augment creativity in such settings.
From research in social psychology, we know that creativity emerges from social interactions in the moment, where people listen, respond, and build on one another (Sawyer, 2019). Like in jazz, one instrument starts to play and others improvise on the spot, creating a completely new tune on the spot. Similarly, in the innovation context, new products might emerge in an exchange of people with different experiences and backgrounds, like designers, computer scientists, and economists. Creativity support design principles emphasize the relevance of exploratory, collaborative, and iterative processes in which ideas are continuously expanded, adapted, and refined (Resnick et al., 2005). However, this iterative and collaborative process remains underexplored in current HCI research on LLMs and creativity.
@@
Psychological theory offers ways to formalize the challenges in human collaborative settings: social loafing refers to the tendency of individuals in a group to contribute less effort compared to when they work alone (Latané et al., 1979; Karau and Williams, 1993). For example, in a brainstorming session, individuals may put less effort into generating novel ideas when working in a large group compared to working alone (Diehl and Stroebe, 1987; Ez-Zaouia and Carrillo, 2024). In the context of suggestions for writing, individuals often prefer the low-effort support of selecting between suggestions compared to actively prompting the suggestions (Dang et al., 2023). Psychological theory suggests that when individuals lack control over an outcome, they may become demotivated and allocate fewer cognitive resources (Latané et al., 1979; Maier and Seligman, 1976). Translated into human–AI collaboration, this tendency can manifest as algorithmic loafing: rather than engaging fully, individuals reduce their own effort and adopt the system’s output with little reflection (Inuwa-Dutse et al., 2023). Applied to our creativity context, this means that, instead of actively striving for the best solution, people may disengage and settle for whatever the AI produces.
@@
As a consequence, in the context of co-creating with LLMs, it might be beneficial to keep humans engaged and in charge of the creative process. Co-creation with LLMs often reduces human effort and makes the task less demanding (Mei et al., 2025). However, research shows that the mode of collaboration affects creative performance. When people co-create with a model rather than merely edit its output, they report higher self-efficacy, which, conversely, leads to higher creative quality (McGuire et al., 2024). This aligns with findings that co-creative AI systems foster higher engagement when users perceive them as collaborative partners that inspire new directions rather than as tools that merely execute commands (Lawton et al., 2023). Consequently, effective co-creation keeps humans involved in the creative process and triggers their creativity rather than allowing mental offloading.
@@
Extending this perspective, creativity research provides support on how to foster creative processes. For example, guidance that prompts designers to actively reflect on their decisions can reduce over-reliance on AI-generated outputs (Gmeiner et al., 2025). Additionally, individuals produce more novel ideas when they are provided analogies from unrelated domains that are still structurally relevant (Dahl and Moreau, 2002). These trigger cognitive processes and allow for producing novel content. CSTs might support here, for example, by providing analogies that are distinct but connected (Gilon et al., 2018). Building on this insight, we argue that human–LLM interaction modes are more effective when they stimulate the human creative process rather than letting the model take the creative lead. For example, when exploring new product ideas, individuals may either become constrained by the model’s output or be inspired by novel analogies they might not have considered otherwise. Here, we distinguish between (1) model-led collaboration, where the LLM generates the core ideas and the human mainly edits or reviews, and (2) human-led collaboration, where the human drives ideation and the LLM supports through providing analogies, useful information, and supporting in elaboration.
Later, we adopt two different human-led interaction modes commonly used in HCI and creativity research: (i) a suggestion mode, where the AI offers alternative ideas, reflects a well-established pattern in creativity support systems that use suggestions to inspire user thinking (Dang et al., 2023; Kumar et al., 2025; Doshi and Hauser, 2024), and (ii) a question mode, where the AI asks reflective questions, draws on work in HCI that employs socratic or reflection-based prompts to encourage users to articulate and extend their ideas (Winkler et al., 2020; Gmeiner et al., 2025; Bastani et al., 2025; Stenning et al., 2016; Al-Hossami et al., 2023). Both interaction modes were designed to stimulate the participants’ thinking and spark their creativity.
@@ 3.2. Group Level: Idea Diversity
While co-creation with LLMs might augment creativity on an individual level, this may still reduce diversity at the group level. Ideas generated by LLMs are more similar to each other compared to ideas generated by a human crowd, implying a possible homogenization effect (Meincke et al., 2025). Also, when humans collaborate with LLMs, the homogenization effect applies, so that ideas (or other text such as stories) may become more similar to each other and thus less diverse (Doshi and Hauser, 2024; Meincke et al., 2024a). A reason for this effect can be human bias, such as fixation effects (Kohn and Smith, 2011; Sio et al., 2015; Crilly, 2019), which also emerge in human-to-human collaborative settings. In contrast to this, we suggest that this effect could be mitigated by interaction modes, where the humans do the main creative work and the model rather acts as a source of inspiration compared to model-led conditions:
@@ 4. Methods
To empirically assess our hypotheses, we performed a preregistered (https://aspredicted.org/nf5r-cmrt.pdf) randomized controlled experiment with $N=502$ participants. Codes and anonymized data are available via our GitHub (https://github.com/SM2982/replication-creativity-CHI2026). Our study compared human-led idea generation (question-mode and suggestion-mode) and model-led idea generation, along with a vanilla LLM condition and a no-LLM control condition. Our study received IRB approval from Ethics Committee at LMU Munich School of Management (ETH-SOM-004, approved March 12, 2025). All participants provided their informed consent and could withdraw at any time. Data were collected and stored in accordance with local data privacy laws. Compensation was done via Prolific, meeting the fair‑wage policy of the platform.
Interfaces for question-mode and suggestion-mode Two interface screenshots with four callouts. Left side shows question-mode as a chat conversation. Callout A highlights a system message that adds domain knowledge about a vehicle feature and asks a reflective question. Callout B highlights a second system message using an analogy (for example, smart home systems) to prompt an alternative idea, followed by the participant response. Right side shows suggestion-mode. Callout C highlights a panel listing three written suggestions for improving novelty and usefulness. Callout D highlights a text box where the participant rewrites their idea based on the suggestions and then finishes.
Refer to caption Interfaces of question-mode and suggestion-mode with example dialogues. The figure shows screenshots of the two human-led conditions. On the left, the question-mode presents a chat-style interface. Panel A shows the system adding knowledge about infrared cameras to inspire alternative ideas. Panel B shows the system providing an analogy to smart home systems to stimulate further creativity. On the right, the suggestion-mode displays three written suggestions that propose ways to expand the initial idea, including intervention, telehealth use, and commercial applications. The text window beneath allows the participant to rewrite and refine the idea based on the suggestions.
Figure 2. Screenshots of the interface of the two human-led conditions. In the question-mode, the LLM provided (A) further knowledge and (B) an analogy to inspire human creativity. After defining three alternative ideas, the user chooses one to elaborate further. In the suggestion-mode, the participants were provided three suggestions on how to make the initial idea more novel, and on how to further elaborate (C). Below the suggestion, a text box is shown where the participants can refine their initial idea based on the LLM input (D).